# Going beyond density functional theory accuracy: Leveraging experimental data to refine pre-trained machine learning interatomic potentials

### v1

@@ Acknowledgments
The authors appreciate the helpful scientific discussions with Nicholas Lubbers, Sakib Matin, Roxanne Tutchton, Michael Cooper, and Kashi Subedi. S.G. is grateful for the mentorship provided by Oleg Prezhdo and Sergei Tretiak. S.G., L.A.V., B.N., A.v.V., T.G., and R.A.M. acknowledge support from Los Alamos National Laboratory (LANL), an affirmative action/equal opportunity employer, operated by Triad National Security LLC, for the National Nuclear Security Administration of the U.S. Department of Energy under Contract No. 89233218CNA000001. The work at LANL was supported by the LANL Laboratory Directed Research and Development (LDRD) Project 20220053DR. S.G. gratefully acknowledges the resources of the Los Alamos National Laboratory (LANL) Computational Science summer student program. This research used resources provided by the Darwin testbed at LANL which is funded by the Computational Systems and Software Environments subprogram of LANLβs Advanced Simulation and Computing program (NNSA/DOE). R.A.M. acknowledges the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.
@@
HIP-NN training works by processing the feature vectors for each atom through interaction blocks consisting of an interaction layers followed by a number of atomic layers. At input layer, the feature vector is simply the one-hot encoding representing the species of an atom (Eg., U, O, N). At subsequent layers, feature vector () of atom has length . At interaction layers, the atom learns from its surrounding atoms within a pre-defined cutoff radius, i.e., the feature vector on one atom mixes with the feature vectors of its neighbouring atoms. Mathematically, the input feature vector is transformed to using:
@@
Here, , and are Cartesian indices that take on values of , , and in some coordinate system, is the Kronecker delta, and is the magnitude squared of the input vector . The interaction term is a weighted sum of rank- tensors:
@@
When UO2 MLIP is trained to the HIP-NN architecture with 1 interaction layer, the RMSE for energy-per-atom and forces decreases for all 3 datasets (train, valid, test) as the number of features increases (TableΒ S1). Interestingly, this is not the case when the number of interaction layers is increased to 2. In this case, while the training dataset RMSE decreases with increasing features, the testing and validation dataset RMSE values (for both energy and forces) are minimum for 60 features instead of 120. The complex neural network architecture with 2 interaction layers and 120 features clearly suffers from overfitting. From this analysis, two combinations of and hyperparameters are selected (interaction layers = 1, number of features = 120) and (interaction layers = 2, number of features = 60). Further, the soft max distance is determined for both sets of hyperparameters (TableΒ S2 and Β S3).
@@
Further we determine the pre-factors of loss function and to train UO2 MLIP with the HIP-NN architecture, by keeping the chosen hyperparameters fixed, (i.e., interaction layers = 1, features = 120, soft max = 4, hard max = 4.7, and soft min = 1.5, atomic layers = 4, sensitivities = 20, lmax=1) in tableΒ S4 below.
From the RMSE values in TableΒ S4, we find that the MLIP trained with the loss function that has higher weight of the force pre-factor results in smaller RMSE values for forces, and MLIP trained with loss function that has higher weight for the energy pre-factor results in smaller RMSE values for energy. Hence, we consider calculating a temperature dependent physical properties such as bulk modulus and compare to experiments to determine the best fit for pre-factors. From FigureΒ S1, we identify that to ratio of 1:100 has bulk modulus closer to experimental values. Moreover, during the re-training, we are fine-tuning the MLIP parameters with only the atomic energies in the refinement dataset, compared to the unrefined MLIP which is pre-trained to both total energy and atomic forces from the DFT dataset. Hence, we considered training the MLIP with relatively higher pre-factors on forces than energy, i.e., =1 and =100.
@@
β’
tensor rank, lmax: 1
@@
β’
tensor rank, lmax: 1
@@
| Unrefined MLIP | 403.9 | 111.1 | 45.7 | 208.4 |
| Refined MLIP | 281.0 | 124.8 | 45.6 | 210.3 |

### v2

@@ Acknowledgments
The authors appreciate the helpful scientific discussions with Nicholas Lubbers, Sakib Matin, Roxanne Tutchton, Michael Cooper, and Kashi Subedi. S.G. is grateful for the mentorship provided by Oleg Prezhdo and Sergei Tretiak. S.G., L.A.V., B.N., A.v.V., T.G., and R.A.M. acknowledge support from Los Alamos National Laboratory (LANL), an affirmative action/equal opportunity employer, operated by Triad National Security LLC, for the National Nuclear Security Administration of the U.S. Department of Energy under Contract No. 89233218CNA000001. R.A.M. acknowledges the Oak Ridge Leadership Computing Facility at the Oak Ridge National Laboratory, which is supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC05-00OR22725.
Funding
The work at LANL was supported by the LANL Laboratory Directed Research and Development (LDRD) Project 20220053DR. S.G. gratefully acknowledges the resources of the Los Alamos National Laboratory (LANL) Computational Science summer student program. This research used resources provided by the Darwin testbed at LANL which is funded by the Computational Systems and Software Environments subprogram of LANL’s Advanced Simulation and Computing program (NNSA/DOE). This research used resources provided by the Los Alamos National Laboratory Institutional Computing Program, which is supported by the U.S. Department of Energy National Nuclear Security Administration under Contract No. 89233218CNA000001.
Author Contributions
R.A.M. conceived and supervised the project. S.G. trained, refined and validated MLIP, and analyzed the results. B.T.N. designed the refinement method and wrote the original refinement code. S.K. introduced and implemented the freezing-layer strategy for transfer learning in the HIP-NN code. S.G. and R.A.M. revised the refinement method and updated the refinement code. L.A.V. supervised the training and molecular dynamics calculations. A.v.V. supervised the EXAFS calculations and provided the experimental EXAFS data. T.G. procured funding and oversaw the project. S.G. and R.A.M wrote the main manuscript draft. L.A.V., B.T.N and S.K. edited and reviewed the manuscript.
Competing interests
The authors declare no competing financial interests.
@@
HIP-NN training works by processing the feature vectors for each atom through interaction blocks consisting of interaction layers followed by a number of atomic layers. At input layer, the feature vector is simply the one-hot encoding representing the species of an atom (Eg., U, O, N). At subsequent layers, feature vector () of atom has length . At interaction layers, the atom learns from its surrounding atoms within a pre-defined cutoff radius, i.e., the feature vector on one atom mixes with the feature vectors of its neighboring atoms. Mathematically, the input feature vector is transformed to using:
@@
Here, , , and are Cartesian indices that take on values of , , and in some coordinate system, is the Kronecker delta, and is the magnitude squared of the input vector . The interaction term is a weighted sum of rank- tensors:
@@
When UO2 MLIP is trained to the HIP-NN architecture with 1 interaction layer, the RMSE for energy-per-atom and forces decreases for all 3 datasets (train, valid, test) as the number of features increases (Table S1). Interestingly, this is not the case when the number of interaction layers is increased to 2. In this case, while the training dataset RMSE decreases with increasing features, the testing and validation dataset RMSE values (for both energy and forces) are minimum for 60 features instead of 120. The complex neural network architecture with 2 interaction layers and 120 features clearly suffers from overfitting. From this analysis, two combinations of and hyperparameters are selected (interaction layers = 1, number of features = 120) and (interaction layers = 2, number of features = 60). Further, the soft max distance is determined for both sets of hyperparameters (Table S2 and S3).
@@
Further we determine the pre-factors of loss function and to train UO2 MLIP with the HIP-NN architecture, by keeping the chosen hyperparameters fixed, (i.e., interaction layers = 1, features = 120, soft max = 4, hard max = 4.7, and soft min = 1.5, atomic layers = 4, sensitivities = 20, =1) in table S4 below.
From the RMSE values in Table S4, we find that the MLIP trained with the loss function that has higher weight of the force pre-factor results in smaller RMSE values for forces, and MLIP trained with loss function that has higher weight for the energy pre-factor results in smaller RMSE values for energy. Hence, we consider calculating a temperature dependent physical property such as bulk modulus and compare to experiments to determine the best fit for pre-factors. From Figure S1, we identify that to ratio of 1:100 has bulk modulus closer to experimental values. Moreover, during the re-training, we are fine-tuning the MLIP parameters with only the atomic energies in the refinement dataset, compared to the unrefined MLIP which is pre-trained to both total energy and atomic forces from the DFT dataset. Hence, we considered training the MLIP with relatively higher pre-factors on forces than energy, i.e., =1 and =100.
@@
tensor rank, : 1
total training parameters: 136456
@@
tensor rank, : 1
total training parameters: 28576
Hyperparameters for training the Aluminum MLIP
number of features: 100
sensitivities: 20
distance, soft min: 1.25
distance, soft max: 7
distance, hard max: 7.5
number of interaction layers: 1
number of atomic layers: 3
sensitivity type: inverse
cusp regularization: 0.001
learning rate: 0.001
tensor rank, : 2
total training parameters: 73356
@@
| Unrefined MLIP | 380.0 | 109.6 | 47.4 | 199.7 |
| Refined MLIP | 370.9 | 121.7 | 46.3 | 204.8 |